OpenStudy (anonymous):

Write the equation of a line with slope m=1/2 and including point (4, -1). y = -2x + 3 y = 2x + 3 1/2 y=x+3 y=1/2 x-3

OpenStudy (matt101):

In y = mx + b form, x and y are the coordinates of your point on the line, m is the slope, and b is the y-intercept. The m and b will be numbers in your general form of the equation, but x and y will stay x and y (because they can represent any point on the line). In this case we have m, x, and y. We can plug the values in to find b: y = mx + b -1 = (1/2)(4) + b b = -3 This means the general form of the equation is: y = (1/2)x - 3 and your answer is D!

OpenStudy (anonymous):

Thank you :)